The graph of y = 1/2x^2 - bx +2 is tangent to the x -axis. Find the value of b.
Answer by Alan3354(69443)

The graph of y = 1/2x^2 - bx +2 is tangent to the x -axis. Find the value of b.
----------
Tangent --> discriminant = 0
b^2 - 4ac = 0
b^2 - 4*(1/2)*2 = 0
b = -2, +2